# Posterior SBC: Simulation-Based Calibration Checking Conditional on Data

### v1

@@ Acknowledgments
We acknowledge the computational resources provided by the Aalto Science-IT project via the Triton compute cluster. TS and AV acknowledge the support by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence, and Research Council of Finland project (340721). MS and PCB acknowledge support of Cyber Valley Project CyVy-RF- 2021-16 and the DFG under Germany’s Excellence Strategy – EXC-2075 - 390740016 (the Stuttgart Cluster of Excellence SimTech). MS acknowledges the European Union’s Horizon 2020 research and innovation programme under grant agreements No 951847 (ELISE) and No 101070617 (ELSA), and the ELLIS PhD program. PCB acknowledges support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526.
Appendix A Case studies: Details and additional results
This supplement provides details and additional results for the case studies of the main paper. The source code for reproducing these case studies is available at https://github.com/TeemuSailynoja/posterior-sbc.
A.1 Case study 2: Lotka-Volterra model
@@
Figure 11: Individual parameter calibration assessments using prior SBC for the Lotka-Volterra model.
Figure 12: Parameter recovery in prior SBC for the Lotka-Volterra model.
In Figure 11 we display the graphical calibration assessment of the individual parameters obtained with prior SBC for the Lotka-Volterra model. Even without correction for multiple comparison, the inference of the parameter posteriors looks well calibrated. In the parameter recovery plots for prior SBC, displayed in Figure 12, we see no systematic biases, but there are some cases where the posterior is very far from the ground truth. As discussed earlier in Section 3.2, these outliers are likely due to the posterior often having distinct modes corresponding to high and low measurement error. This leads to a high probability of one or more of the MCMC chains not properly exploring the full posterior.
Figure 13 and Figure 14 show the calibration assessment and recovery of the individual parameters once we condition the inference on the historical observations. Again, we observe no calibration issues. The parameter recovery also exhibits no outliers similar to those observed in prior SBC.
Figure 13: Individual parameter calibration assessments using posterior SBC for the Lotka-Volterra model.
Figure 14: Parameter recovery in posterior SBC for the Lotka-Volterra model.
A.2 Case study 3: Posterior SBC checking in amortized Bayesian inference
@@
Figure 15 shows the parameter recovery of the amortized approximator across a range of settings. We observe that most parameters can be recovered with sufficiently high accuracy given the number of available observations in each dataset (see 15(a)). Further, the epistemic uncertainty in the parameter estimates shrinks as the number of observations grows (see 15(b)). Finally, errors in the posterior estimates propagate into the parameter recovery based on simulated data from the approximate posterior predictive distribution (see 15(c),15(d)). This evaluation follows the same principle as posterior SBC but reports the data-conditional parameter recovery rather than the data-conditional simulation-based calibration results. In summary, the patterns from the empirical parameter recovery match the results from (prior and posterior based) simulation-based calibration checking in Section 3.3.
@@
Paralleling the previous evaluation for model $M_{1}$, Figure 16 reports the parameter recovery capabilities of the amortized approximator. Again, most parameters can be recovered with sufficiently high accuracy given the number of available observations in each dataset (see 16(a)), and the uncertainty shrinks when the datasets contain more observations (see 16(b)). The large epistemic uncertainty in the estimate of the boundary collapse parameter $\lambda$ has previously been reported by Ghaderi-Kangavari et al., (2023). Furthermore, the parameter recovery patterns mirror the results from (prior and posterior based) simulation-based calibration checking in Section 3.3. Most notably, the previously reported data-conditional miscalibration in the parameter $\mu_{(e)}$ (10(b)) manifests in a biased data-conditional parameter recovery in 16(c),16(d).

### v2

@@ Acknowledgments
We thank Martin Modrák, Valentin Pratz, and Lasse Elsemüller for helpful comments and suggestions, and we acknowledge the computational resources provided by the Aalto Science-IT project via the Triton compute cluster. TS and AV acknowledge the support by the Research Council of Finland Flagship programme: Finnish Center for Artificial Intelligence, and Research Council of Finland project (340721). MS and PCB acknowledge support of Cyber Valley Project CyVy-RF- 2021-16 and the DFG under Germany’s Excellence Strategy – EXC-2075 - 390740016 (the Stuttgart Cluster of Excellence SimTech). MS acknowledges the European Union’s Horizon 2020 research and innovation programme under grant agreements No 951847 (ELISE) and No 101070617 (ELSA), and the ELLIS PhD program. PCB acknowledges support of the DFG Collaborative Research Center 391 (Spatio-Temporal Statistics for the Transition of Energy and Transport) – 520388526.
Appendix A: Case study 2: Lotka-Volterra model
@@
(a) Individual parameter calibration assessments.
(b) Parameter recovery.
Figure 11: Per parameter results of prior SBC for the Lotka-Volterra model.
In Figure 11(a) we display the graphical calibration assessment of the individual parameters obtained with prior SBC for the Lotka-Volterra model. Even without correction for multiple comparison, the inference of the parameter posteriors looks well calibrated. In the parameter recovery plots for prior SBC, displayed in Figure 11(b), we see no systematic biases, but there are some cases where the posterior is very far from the ground truth. As discussed in Section 3.2 of the article, these outliers are likely due to the posterior having distinct modes, which leads to a high probability of one or more MCMC chains not properly exploring the full posterior.
Figure 12(a) and Figure 12(b) show the calibration assessment and recovery of the individual parameters once we condition the inference on the historical observations. Again, we observe no calibration issues. The parameter recovery also exhibits no outliers similar to those observed in prior SBC.
(a) Individual parameter calibration assessments.
(b) Parameter recovery.
Figure 12: Per parameter results of posterior SBC for the Lotka-Volterra model.
Appendix B: Case study 3: Posterior SBC checking in amortized Bayesian inference
@@
Figure 13 shows the parameter recovery of the amortized approximator across a range of settings. We observe that most parameters can be recovered with sufficiently high accuracy given the number of available observations in each dataset (see Figure 13(a)). Further, the epistemic uncertainty in the parameter estimates shrinks as the number of observations grows (see Figure 13(b)). Finally, errors in the posterior estimates propagate into the parameter recovery based on simulated data from the approximate posterior predictive distribution (see Figure 13(c),13(d)). This evaluation follows the same principle as posterior SBC but reports the data-conditional parameter recovery rather than the data-conditional simulation-based calibration results. In summary, the patterns from the empirical parameter recovery match the results from (prior and posterior based) simulation-based calibration checking in Section 2.3.
@@
Paralleling the previous evaluation for model $M_{1}$, Figure 14 reports the parameter recovery capabilities of the amortized approximator. Again, most parameters can be recovered with sufficiently high accuracy given the number of available observations in each dataset (see Figure 14(a)), and the uncertainty shrinks when the datasets contain more observations (see Figure 14(b)). The large epistemic uncertainty in the estimate of the boundary collapse parameter $\lambda$ has previously been reported by Ghaderi-Kangavari et al., (2023). Furthermore, the parameter recovery patterns mirror the results from (prior and posterior based) simulation-based calibration checking in Section 3.3. Most notably, the previously reported data-conditional miscalibration in the parameter $\mu_{(e)}$ (Figure 10b in the article) manifests in a biased data-conditional parameter recovery in Figure 14(c),14(d).